write an equation for the line, in point slope form, that passes through the points (-5,7) and (-4,-3)
Answer by venugopalramana(3286)

write an equation for the line, in point slope form, that passes through the points (-5,7) and (-4,-3)
formula for eqn.of line connecting 2 points x1,y1 and x2,y2 is given by
y-y1=(x-x1)*{(y2-y1)/(x2-x1)}....and slope of the line =(y2-y1)/(x2-x1)
so eqn.of line is y-(-3)=(x-(-4))*{(-3-(7))/(-4-(-5))
y+3=(x+4)*(-10/1)=-10(x+4)=-10x-40
y=-10x-40-3
y=-10x-43